# Multidimensional Dickman distribution and operator selfdecomposability

### v1

@@ 4 Some applications of the operator Dickman distributions
where $n\geq 1$ and $L(0)\coloneq 0$. Let us consider, for example, $p_{k}=\alpha/k,k\geq 1$ for some $\alpha\in(0,1)$. It then holds that
@@
$=\left(1-\frac{\alpha}{m+1}\right)\cdot\ldots\cdot\left(1-\frac{\alpha}{m+r-1}\right)\frac{\alpha}{m+r},$
@@
We denote $\bar{\mathbb{R}}^{d}$ a one-point compactification of $\mathbb{R}^{d}$ and $\bar{\mathbb{R}}^{d}_{0}=\bar{\mathbb{R}}^{d}\setminus\{0\}$, $\bar{\mathbb{R}}_{+}=(0,\infty]$. We also introduce for any $Q\in\mathscr{M}_{+}$ the set $S_{Q}$ given by
@@
Recall now that for $Q\in\mathscr{M}_{+}$ and $\sigma\in\mathscr{S}$ the Lévy measure of the distribution $\mathscr{D}(Q,\sigma)$ is given by
@@
$\hat{\varrho}(\mathbf{z})=\exp\left[\int\limits_{\mathbb{R}^{d}}(e^{i\mathbf{z}\cdot\mathbf{x}}-1)\mu(d\mathbf{x})\right].$ (22)
@@
Fix $Q\in\mathscr{M}_{+}$ and $\sigma\in\mathscr{S}$. Let $\varrho_{1},\varrho_{2},\ldots\in ID(\mathbb{R}^{d})$ be a sequence of infinitely divisible distributions satisfying the conditions (21) and (22). Let $\mu_{1},\mu_{2},\ldots$ denote the sequence of their Lévy measures.
@@
$\mu_{n}(A_{t,D})\to\log(1\wedge t)\sigma(D),\;n\to\infty,$ (23)
@@
Due to [KEN99, Th. 8.7] (see also [KAL21, Th. 7.7]) it is enough to check the following three conditions
@@
    $\lim\limits_{\delta\downarrow 0}\limsup\limits_{n\to\infty}\int\limits_{|\mathbf{x}|\leq\delta}(\mathbf{z}\cdot\mathbf{x})^{2}\mu_{n}(d\mathbf{x})=0$ for any $\mathbf{z}\in\mathbb{R}^{d}$;
@@
It follows easily from Lemma 4.3 and (20) that conditions (i) and (23) are equivalent in the case of the $\mathscr{D}(Q,\sigma)$ distribution. Notice next that
$\lim\limits_{n\to\infty}\int\limits_{|\mathbf{x}|\leq 1}|\mathbf{x}|\mu_{n}(d\mathbf{x})\leq\lim\limits_{n\to\infty}\int\limits_{\mathbb{R}^{d}}(|\mathbf{x}|\wedge 1)\mu_{n}(d\mathbf{x})=\int\limits_{\mathbb{R}^{d}}(|\mathbf{x}|\wedge 1)M(d\mathbf{x})<\infty,$
where we used the definition of vague convergence in the second equality and Corollary 3.6 on the last step. Thus, assuming that (i) holds, we can easily obtain (ii). Indeed, we have
$\lim\limits_{\delta\downarrow 0}\limsup\limits_{n\to\infty}\int\limits_{|\mathbf{x}|\leq\delta}(\mathbf{z}\cdot\mathbf{x})^{2}\mu_{n}(d\mathbf{x})\leq|\mathbf{z}|^{2}\lim\limits_{\delta\downarrow 0}\delta\cdot\int\limits_{|\mathbf{x}|\leq 1}|\mathbf{x}|M(d\mathbf{x})=0.$
@@
Since for each $k=1,\ldots,d$ the component $x_{k}c(\mathbf{x})$ defines the bounded and continuous map $\mathbb{R}^{d}\mapsto\mathbb{R}$, which vanishes on a neighbourhood of $\mathbf{0}$ (see, in particular, conditions [KEN99, (8.3), (8.4)]), the condition (iii) follows directly from (i) and the definition of vague convergence. ∎
@@
The next corollary provides a tractable example of the sequence $\mu_{n}$ (and therefore corresponding $ID$ distributions) satisfying the conditions of Theorem 4.4. It is a direct generalisation of [GZ24a, Cor. 1].
@@
Let $\varrho_{\varepsilon}$, $\varepsilon>0$ be a collection of $ID(\mathbb{R}^{d})$ measures satisfying (22) and assume that the corresponding Lévy measures $\mu_{\varepsilon}$, $\varepsilon>0$, are of the form
$\mu_{\varepsilon}(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{1}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\varepsilon\rho(\varepsilon r,\mathbf{x})\sigma(d\mathbf{x})dr,\quad B\in\mathscr{B}(\mathbb{R}^{d}).$
Assume also that for some $h:\mathbb{S}^{d-1}\mapsto\mathbb{R}$, $h\geq 0$, it holds that
@@
$\varrho_{\varepsilon}\overset{w}{\to}\mathscr{D}(Q,\tilde{\sigma})\;\text{ as }\varepsilon\downarrow 0,$
where $\tilde{\sigma}(B)=\int\limits_{B}h(\mathbf{x})\sigma(d\mathbf{x}),B\in\mathscr{B}(\mathbb{S}^{d-1})$.
Remark 10.
Let us consider the $ID(\mathbb{R}^{d})$ measure $\varrho$ satisfying (22) with Lévy measure $\tilde{\mu}$ given as
$\tilde{\mu}(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\infty}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\rho(r,\mathbf{x})\sigma(d\mathbf{x})dr,\quad B\in\mathscr{B}(\mathbb{R}^{d}).$
For the measure $\tilde{\mu}$ and some $u>0$ we consider the truncation
$T_{u}\tilde{\mu}(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{u}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\rho(r,\mathbf{x})\sigma(d\mathbf{x})dr,\quad B\in\mathscr{B}(\mathbb{R}^{d}),$
and note that $T_{u}\tilde{\mu}$ is a Lévy measure for any $u>0$. Note also that the truncation $T_{u}$ corresponds to considering only small jumps of the underlying Poisson random measure, whereas $\tilde{\mu}-T_{u}\tilde{\mu}$ corresponds to the compound Poisson part. Denote then $T_{u}\varrho$ the $ID(\mathbb{R}^{d})$ distribution with Lévy measure $T_{u}\tilde{\mu}$. Then $\varrho_{\varepsilon}$ of Corollary 4.5 can be obtained by the transformation $\varrho\mapsto\varepsilon^{-Q}T_{\varepsilon}\varrho$, which accords with truncation and operator rescaling of the original measure. Similar constructions have been considered in more detail in [COV09] and [GZ24a]. The application of this result for simulations has also been discussed there.
@@
The proof is analogous to the proof of 3.8iii if we notice that
$\frac{c_{1}+\cdots+c_{k}}{c}\frac{c_{1}\nu_{1}(\cdot)+\cdots+c_{k}\nu_{k}(\cdot)}{c_{1}+\cdots+c_{k}}\overset{v}{\to}\tilde{\nu}\;\text{as}\;k\to\infty$
by Lemma 4.3 and use Theorem 4.4. ∎

### v2

@@ 4 Some applications of the operator Dickman distributions
where $n\geq 1$ and $L(0)\coloneq 0$. Let us consider, for example, $p_{k}=\alpha/k$, $k\geq 1$, for some $\alpha\in(0,1)$. It then holds that
@@
$=\left(1-\frac{\alpha}{m+1}\right)\cdot\ldots\cdot\left(1-\frac{\alpha}{m+r-1}\right)\frac{\alpha}{m+r},\;r\geq 1$
@@
We denote $\bar{\mathbb{R}}^{d}$, a one-point compactification of $\mathbb{R}^{d}$ and $\bar{\mathbb{R}}^{d}_{0}=\bar{\mathbb{R}}^{d}\setminus\{0\}$, $\bar{\mathbb{R}}_{+}=(0,\infty]$. We also introduce for any $Q\in\mathscr{M}_{+}$ the set $S_{Q}$ given by
@@
Recall now that for $Q\in\mathscr{M}_{+}$ and $\sigma\in\mathscr{S}$, the Lévy measure of the distribution $\mathscr{D}(Q,\sigma)$ is given by
@@
$\hat{\varrho}(\mathbf{z})=\exp\left[\int_{\mathbb{R}^{d}}(e^{i\mathbf{z}\cdot\mathbf{x}}-1)\mu(d\mathbf{x})\right].$ (23)
@@
Fix $Q\in\mathscr{M}_{+}$ and $\sigma\in\mathscr{S}$. Let $\varrho_{1},\varrho_{2},\ldots\in ID(\mathbb{R}^{d})$ be a sequence of infinitely divisible distributions and let $\mu_{1},\mu_{2},\ldots$ denote the sequence of their Lévy measures, which we assume satisfy the condition (22). Assume additionally that
$\lim_{\delta\to 0}\limsup_{n\to\infty}\int_{|\mathbf{x}|<\delta}|\mathbf{x}|\mu_{n}(d\mathbf{x})=0.$ (24)
@@
$\mu_{n}(A_{t,D})\to-\log(1\wedge t)\sigma(D),\;n\to\infty,$ (25)
@@
Due to [KEN99, Th. 8.7], it is enough to check the following three conditions
@@
    $\lim\limits_{\delta\to 0}\limsup\limits_{n\to\infty}\int\limits_{|\mathbf{x}|\leq\delta}(\mathbf{z}\cdot\mathbf{x})^{2}\mu_{n}(d\mathbf{x})=0$ for any $\mathbf{z}\in\mathbb{R}^{d}$;
@@
It follows from Lemma 4.3 and (21) that conditions (i) and (25) are equivalent in the case of the $\mathscr{D}(Q,\sigma)$ distribution. Notice also that the condition (ii) follows readily from (24). Consider condition (iii) now. Assuming that (i) holds, we can rewrite for any $\delta>0$
$\lim\limits_{n\to\infty}\int\limits_{\mathbb{R}^{d}}\mathbf{x}c(\mathbf{x})\mu_{n}(d\mathbf{x})=$ $\lim\limits_{n\to\infty}\left(\int\limits_{|\mathbf{x}|<\delta}\mathbf{x}c(\mathbf{x})\mu_{n}(d\mathbf{x})+\int\limits_{|\mathbf{x}|\geq\delta}\mathbf{x}c(\mathbf{x})\mu_{n}(d\mathbf{x})\right)$
$=$ $\lim\limits_{n\to\infty}\int\limits_{|\mathbf{x}|<\delta}\mathbf{x}c(\mathbf{x})\mu_{n}(d\mathbf{x})+\int\limits_{|\mathbf{x}|\geq\delta}\mathbf{x}c(\mathbf{x})M(d\mathbf{x}),$
@@
where the last equality follows from the definition of vague convergence since the maps $\mathbb{R}^{d}\mapsto\mathbb{R}$, $\mathbf{x}\mapsto x_{k}c(\mathbf{x})\mathds{1}{|\mathbf{x}|\geq\delta}$, $k=1,\ldots,d$, define bounded continuous functions, which vanish at the neighbourhood of $\mathbf{0}$. We can now let $\delta\to 0$ and (iii) follows from (24). ∎
We now consider the specific example of a sequence satisfying the conditions of Theorem 4.4. Namely let $\mu$ be a Lévy measure on $\mathbb{R}^{d}$ such that
$\mu(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\infty}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\varpi(d\mathbf{x},dr),\quad B\in\mathscr{B}(\mathbb{R}^{d}),$ (26)
where $Q\in\mathscr{M}_{+}$ is fixed and $\varpi(\cdot,\cdot)$ is a measure on $\mathbb{S}^{d-1}\times\mathbb{R}_{+}$ such that
$\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\infty}|r^{Q}\mathbf{x}|\mathds{1}_{B_{1}}(r^{Q}\mathbf{x})\varpi(d\mathbf{x},dr)<\infty,$ (27)
where we denote $B_{h}\subset\mathbb{R}^{d}$, an open ball of radius $h>0$. Thus, measure $\mu$ satisfies (22) and there exists $\rho\in ID(\mathbb{R}^{d})$ such that (23) holds.
For the measure $\mu$ and some $u>0$, we consider the truncation
$T_{u}\mu(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{u}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\varpi(d\mathbf{x},dr),\quad B\in\mathscr{B}(\mathbb{R}^{d}),$
and note that $T_{u}\mu$ is a Lévy measure for any $u>0$. Note also that the truncation $T_{u}$ corresponds to considering only small jumps of the underlying Poisson random measure, whereas the measure $\mu-T_{u}\mu$ corresponds to the compound Poisson part. Denote then $T_{u}\varrho$, the $ID(\mathbb{R}^{d})$ distribution with Lévy measure $T_{u}\mu$. Consider further the transformation $\varrho\mapsto\varepsilon^{-Q}T_{\varepsilon}\varrho$, $\varepsilon>0$, which accords with truncation and operator rescaling of the original measure and denote $\varrho_{\varepsilon}=\varepsilon^{-Q}T_{\varepsilon}\varrho$. Let $\mu_{\varepsilon}$ be the corresponding Lévy measure, i.e.
$\mu_{\varepsilon}(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\varepsilon}\mathds{1}_{B}\{(r/\varepsilon)^{Q}\mathbf{x}\}\varpi(d\mathbf{x},dr),\quad B\in\mathscr{B}(\mathbb{R}^{d}).$ (28)
We show further that for measures $\mu_{\varepsilon}$, $\varepsilon\to 0$, the condition (24) follows from (25). Consequently, the vague convergence of the Lévy measures $\mu_{\varepsilon}\overset{v}{\to}M$ is sufficient to establish weak convergence $\varrho_{\varepsilon}\overset{w}{\to}\mathscr{D}(Q,\sigma)$, $\varepsilon\to 0$, for a sequence $\varrho_{\varepsilon}$ obtained by the transformation of measure $\varrho$. We note that similar constructions have been considered in more detail in [COV09] and [GZ24a]. The application of this result for simulations has also been discussed there.
Lemma 4.5.
Let $\mu_{\varepsilon}$, $\varepsilon>0$, be a collection of Lévy measures given by (28). Assume also that there exists some probability measure $\sigma$ on $\mathbb{S}^{d-1}$ such that for every $t>0$ and every $D\in\mathscr{B}(S_{Q})$, $\sigma(\partial D)=0$, it holds that
$\mu_{\varepsilon}(A_{t,D})\to-\log(1\wedge t)\sigma(D),\;\varepsilon\to 0,$
where $A_{t,D}$ is given as in Lemma 4.3. Then $\varrho_{\varepsilon}\overset{w}{\to}\mathscr{D}(Q,\sigma)$.
Proof.
Note the assumption above implies that $\mu_{\varepsilon}\overset{v}{\to}M$, where measure $M$ is given as in (21), see Lemma 4.3. We will now show that
$\lim_{\delta\to 0}\limsup_{\varepsilon\to 0}\int_{|\mathbf{x}|<\delta}|\mathbf{x}|\mu_{\varepsilon}(d\mathbf{x})=0.$ (29)
The idea of the proof is due to [GZ24a, Prop. 1]. Rewrite now for some $\delta,\varepsilon>0$
$\int_{|\mathbf{x}|<\delta}|\mathbf{x}|\mu_{\varepsilon}(d\mathbf{x})=$ $\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\varepsilon}|(r/\varepsilon)^{Q}\mathbf{x}|\mathds{1}_{B_{\delta}}((r/\varepsilon)^{Q}\mathbf{x})\varpi(d\mathbf{x},dr)$
$=$ $\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{1}|r^{Q}\mathbf{x}|\mathds{1}_{B_{\delta}}(r^{Q}\mathbf{x})\varpi(d\mathbf{x},\varepsilon dr).$
Using (11), we get $|r^{Q}\mathbf{x}|\geq c_{1}r^{K}|\mathbf{x}|=c_{1}r^{K}$ for all $0\leq r\leq 1$ and some $c_{1},K>0$. Denote then $g(\delta)=(\delta/c_{1})^{1/K}$ and notice that
$\int_{|\mathbf{x}|<\delta}|\mathbf{x}|\mu_{\varepsilon}(d\mathbf{x})\leq$ $\int_{\mathbb{S}^{d-1}}\int_{0}^{g(\delta)}|r^{Q}\mathbf{x}|\varpi(d\mathbf{x},\varepsilon dr)$
$=$ $\sum_{n=1}^{\infty}\int_{\mathbb{S}^{d-1}}\int_{g(\delta)2^{-n}}^{g(\delta)2^{-n+1}}|r^{Q}\mathbf{x}|\varpi(d\mathbf{x},\varepsilon dr)$
$=$ $\sum_{n=1}^{\infty}\int_{\mathbb{S}^{d-1}}\int_{1/2}^{1}\left|\left(\frac{rg(\delta)}{2^{n-1}}\right)^{Q}\mathbf{x}\right|\varpi\left(d\mathbf{x},\varepsilon\frac{g(\delta)}{2^{n-1}}dr\right)$
$\leq$ $\sum_{n=1}^{\infty}c_{2}\left(\frac{g(\delta)}{2^{n-1}}\right)^{k}\int_{\mathbb{S}^{d-1}}\int_{1/2}^{1}\left|r^{Q}\mathbf{x}\right|\varpi\left(d\mathbf{x},\varepsilon\frac{g(\delta)}{2^{n-1}}dr\right),$
where we applied (11) with $s=g(\delta)/2^{n-1}<1$, assuming that $\delta$ is small enough. Notice now that
$\int_{\mathbb{S}^{d-1}}\int_{1/2}^{1}\left|r^{Q}\mathbf{x}\right|\varpi\left(d\mathbf{x},\varepsilon\frac{g(\delta)}{2^{n-1}}dr\right)=\int_{\mathbb{R}^{d}}|\mathbf{y}|\mathds{1}_{A_{1/2}^{1}}(\mathbf{y})\mu_{\varepsilon\frac{g(\delta)}{2^{n-1}}}(d\mathbf{y}),$
where $A_{a}^{b}=\{\mathbf{y}\in\mathbb{R}^{d}\,|\mathbf{y}=r^{Q}\mathbf{x},r\in(a,b),\mathbf{x}\in\mathbb{S}^{d-1}\}$, $0\leq a<b<\infty$. By the definition of vague convergence, the last integral converges as $\varepsilon\to 0$ to
$\int_{\mathbb{R}^{d}}|\mathbf{y}|\mathds{1}_{A_{1/2}^{1}}(\mathbf{y})M(d\mathbf{y})=m^{*}<\infty.$
Consequently, for any $\zeta>0$ there exists $\varepsilon^{*}>0$ such that for every $0<\varepsilon<\varepsilon^{*}$, it holds that
$\int_{\mathbb{R}^{d}}|\mathbf{y}|\mathds{1}_{A_{1/2}^{1}}(\mathbf{y})\mu_{\varepsilon\frac{g(\delta)}{2^{n-1}}}(d\mathbf{y})<m^{*}+\zeta.$
The last estimate holds for all $n\geq 1$ and for $\delta$ small enough; indeed, since function $g(\delta)$ decreases to 0 monotonically as $\delta\to 0$ it follows that $\varepsilon\frac{g(\delta)}{2^{n-1}}<\varepsilon<\varepsilon^{*}$ for $\delta$ small. Consequently, we can deduce that for $\delta$ small, it holds that
$\int_{|\mathbf{x}|<\delta}|\mathbf{x}|\mu_{\varepsilon}(d\mathbf{x})\leq c_{2}(g(\delta))^{k}\sum_{n=1}^{\infty}2^{-k(n-1)}(m^{*}+\zeta).$
We can now conclude that (29) holds, which ends the proof by applying Theorem 4.4. ∎
@@
The next corollary provides a tractable example of the possible measure $\mu$ (and therefore corresponding $ID$ distribution) satisfying the conditions of Lemma 4.5. It is a direct generalisation of [GZ24a, Cor. 1].
@@
Consider $\varrho\in ID(\mathbb{R}^{d})$ and assume that its Lévy measure $\mu$, satisfying (27), can be given by
$\mu(B)=\int\limits_{\mathbb{S}^{d-1}}\int\limits_{0}^{\infty}\mathds{1}_{B}\{r^{Q}\mathbf{x}\}\rho(\mathbf{x},r)\sigma(d\mathbf{x})dr,\quad B\in\mathscr{B}(\mathbb{R}^{d}),$
i.e. assume that $\varpi(d\mathbf{x},dr)=\rho(\mathbf{x},r)\sigma(d\mathbf{x})dr$ in (26) for some function $\rho:\mathbb{S}^{d-1}\times(0,\infty)\mapsto[0,\infty)$ and a sigma-finite measure $\sigma$ on $\mathbb{S}^{d-1}$.
Assume also that for some $h:\mathbb{S}^{d-1}\mapsto\mathbb{R}_{+}$, it holds that
@@
$\varrho_{\varepsilon}\overset{w}{\to}\mathscr{D}(Q,\tilde{\sigma})\;\text{ as }\varepsilon\to 0,$
where $\tilde{\sigma}(B)=\int\limits_{B}h(\mathbf{x})\sigma(d\mathbf{x})$, $B\in\mathscr{B}(\mathbb{S}^{d-1})$, and $\varrho_{\varepsilon}$ is obtained from $\varrho$ by the transformation described above.
@@
The proof is analogous to the proof of Proposition 3.8iii if we notice that
$\frac{w_{1}\nu_{1}+\cdots+w_{k}\nu_{k}}{w}(D)\begin{cases}\leq\tilde{\nu}(D),&k\in\mathbb{N},\\ \to\tilde{\nu}(D)&\text{as}\;k\to\infty.\end{cases}$
for all $D\subset\mathbb{S}^{d-1}$, and use Theorem 4.4. ∎